# Protein Structure Tokenization via Geometric Byte Pair Encoding

### v1

@@ 3.3 Principles of Protein Structure Tokenization
Principle 3: Downstream transfer via codebook/vocabulary. Let $\mathcal{V}$ be the vocabulary of $\mathsf{T}$ and let $N(\mathbf{x})$ be the residue count. Let $\Theta$ parameterize a pretrained feature extractor. The codebook/vocabulary $\mathcal{V}$ induces per-residue features $r_{\mathcal{V}}(\mathbf{x})=\Psi_{\mathcal{V}}(F_{\Theta}(\mathbf{x}))\in(\mathbb{R}^{d})^{N(\mathbf{x})}$. An ideal tokenizer of protein structures should go beyond pure compression; it should learn useful signals related to function. We loosely define the ability of a PST to transfer useful signals by test performance on a battery of downstream tasks when parameterizing samples $\mathbf{x}$ by the vocabulary $\mathcal{V}$ together with a feature extractor $\Theta$. We benchmark downstream transfer of GeoBPE against others in Table 1.
Figure 2: (Top) GeoBPE tracks a Geo-Pair Encoding, a dictionary mapping Geo-Pair keys to occurrences at all times. Each step pops the most frequent Geo-Pair key, gathers the occurrences and selects $K$ prototypes to add to $\mathcal{V}$. All occurrences are quantized to the closest prototype. Glue angles are optimized to correct for the drift introduced. (Bottom) Toy example with two backbones; we initialize residue-orientation modes using two prototypes (green & yellow), pop the frequent Geo-Pair (blue), quantize occurrences (rounded $\rightarrow$ sharp corners), and optimize glue angles (red and orange).
@@ 4 Experiments
  • Token Efficiency: (Q5) How frequent and balanced is vocabulary utilization? (Q6) Does small-scale autoregressive token modeling generate reasonable designs?
@@
Downstream Transfer. For VQ-VAEs, $\Theta$ and $\mathcal{V}$ are jointly learned, so we set $r^{\text{VQ-VAE}}_{\mathcal{V}}(\textbf{x}):=r^{\text{VQ-VAE}}(\textbf{x})\leftarrow\text{Enc}(\textbf{x})$. For GeoBPE, we use $\Theta\leftarrow\text{ESM3}$ to demonstrate how $\mathcal{V}^{\text{{GeoBPE}}}$ can transfer useful signals from $\mathcal{F}_{\Theta}(\textbf{x})$ to $r_{\mathcal{V}}(\mathcal{F}_{\Theta}(\textbf{x}))$.
Performance Metrics. Compression measures Bits-Per-Residue (BPR), as defined in Sec. 3.3. Distortion ($\Delta$) use standard RMSD and LDDT. Token Efficiency uses Codebook Utility Rate (UR), Perplexity (details in App. D) and Small Structure Language Model Evaluation (SSLM-Eval) (details in App. E). SSLM-Eval trains a small decoder-only Transformer on the same pretraining corpus post-tokenization (Algo. 2), then generates new sentences, detokenizes them into structures, and assesses generation results (Algo. 3, 4, 5). Downstream Transfer covers 12 tasks (24 test splits) using AUROC (%) for functional site prediction, Spearman’s $\rho$ (%) for flexibility prediction, and Macro F1 (%) for fold classification. Expert Agreement measures Domain & Segment Recall/Precision/F1/IOU (details in App. F).
Computational Complexity / Implementation Details. We analyze the theoretical complexity of GeoBPE in App. H and justify the steps we took towards efficient implementation and use.
@@ 5 Results
Figure 3: Plots of $(\operatorname{BPR}(\mathsf{T};\mathcal{D}),\ \Delta(\mathsf{T};\mathcal{D}_{\mathrm{test}}))$ across tokenizers for $\Delta\in\{\text{RMSD},\text{LDDT}\}$. We vary $|\mathcal{V}|\in\{128,256,512,1024\}$ for VQ-VAE/AminoASeed and $|\mathcal{V}|\in\{600,2500,6000,21000\}$ for GeoBPE to sample multiple points; we observe GeoBPE sweeps a smooth tradeoff curve.
@@
Token Efficiency. We report UR & Perplexity averaged over held-out test sets to gauge codebook/vocabulary usage on unseen data, the setting where the tokenizer is deployed. In Table 6, we see all methods except VQ-VAE and ESM3 achieve an average UR of $40\%$; all except VQ-VAE achieve $0.2$ average Perplexity. An ideal tokenizer avoids codebook collapse, but exactly uniform token usage may not be desirable. We introduce SSLM-Eval to stress test whether codebook efficiency actually leads to generative efficiency. SSLM-Eval is a holistic way to compare tokenizers using both encoder token efficiency and decoder’s generative efficiency. In Table 7, we find GeoBPE is capable of generating $99\%$ unique and designable backbones, achieving up to $49\%$ higher scTM and maintaining higher diversity than both VQ-VAE methods using the same data splits. We visualize some realistic, novel backbones GeoBPE generated in App. E.4. Interestingly, the “less-efficient” VQ-VAE generated $58\%$ more diverse backbones, demonstrating uniform token usage can be counterproductive to language modeling.
Downstream Task Transfer. In Table 1, we see GeoBPE-induced features rank first, on average, across both function and structure property prediction tasks. The relative performance gaps $15.44\%$ and $43.28\%$ quantify the add-on benefits of GeoBPE-induced features. GeoBPE-induced features reverse the trend that discrete PSTs produce less informative representations for downstream tasks (due to quantization-related issues (Yuan et al., 2025)), highlighting that hierarchical structure from discrete vocabularies raises the ceiling on downstream transfer. We include a series of further ablation studies in App. A, where key findings include (1) GeoBPE shows no drop in performance when using only 1% of the pretraining data (all hyperparameters held constant); (2) GeoBPE task-specific oracle tokenizers also show no drop in performance.
Table 1: Downstream transfer performance benchmark. The relative performance v.s. ESM3 for GeoBPE is included. We underline and bold the best continuous and discrete PSTs, respectively;   indicates the best method across both continuous and discrete PSTs. Omitted rows are in Table 4.
@@
ProteinMPNN MIF FoldSeek ProTokens ESM3 VanillaVQ AminoAseed GeoBPE (v.s. ESM3)
@@
[…]
SupFam $0 \ul94$.56 46.26 86.05 90.77 86.71 90.53 91.31 (+0.59%)
[…]
@@ 6 Discussion
Table 2: We annotate 100 PDBs from each dataset and report % of 1,000 random equal-length segmentations that GeoBPE matches or outscores. Omitted columns are in Table 5.
@@
We ran CATH Functional Families (FunFams) (Das et al., 2015b) to obtain domain boundaries and compared them against GeoBPE-derived motifs. Because sequence conservation is linked to structural preservation, we expect overlap between predicted motifs and functional domains. In Table 2, GeoBPE achieves 99.97% domain recall with mean F1 = 0.996 and IOU = 0.992, showing near-perfect agreement across 10 datasets. The agreement is not only geometric but also functional: GeoBPE tokens frequently coincide with boundaries of ligand-binding grooves, transmembrane cavities, and scaffolding helices, capturing motifs that underlie molecular recognition and catalysis. This suggests GeoBPE does more than segment folds consistently: it surfaces interpretable structural primitives that map onto biochemical roles, offering a functional vocabulary absent in prior PSTs. Details are in App. F
Case Study: Human Expert Analysis of Interpretability. We conducted three expert evaluations of GeoBPE-derived hierarchies (App. G). Across proteins, the discovered motifs align with functionally meaningful substructures, including regions mediating ligand binding, molecular recognition, and structural gating. In the SLC25A20 transporter (Fig. 8), GeoBPE isolates a transmembrane binding cavity formed by helices and polar residues. In the 14-3-3:Tau complex (Fig. 9), it identifies a canonical phospho-binding groove stabilized by charged side chains. Recurrent local motifs (aromatic cages, polar bridges, helix-helix clamps) are combined into higher-order scaffolds that mirror established biochemical organization. These hierarchies capture geometric regularities and also modular design principles conserved across folds and families. Even in compact domains, such as nucleotide-recognition modules, GeoBPE motifs reveal the coupling between geometric curvature and chemical specificity, meaning that GeoBPE surfaces reusable motifs that are both interpretable and evolutionarily grounded.
Limitations. GeoBPE currently does not incorporate sequence or side chains, and $\mathcal{V}$-induced features remain dependent on pretrained PLMs. While we tested vocabularies up to 21K tokens, scalability to even larger codebooks and the associated computational tradeoffs remain open questions. The present generative integration is limited to a small-scale proof of concept, and extending it to more powerful models will require further development. Future work should also investigate incorporating side-chain geometry to capture finer-grained biochemical interactions.
@@ 7 Conclusion
8 Reproducibility Statement
We have included detailed descriptions of our method in the main text, to the extent we believe is sufficient to reproduce our method. In App. K, we include mathematical descriptions and algorithm pseudocodes for all mentioned algorithms and subroutines of a non-trivial nature. In App. I, we list the key hyperparameters, their effects on algorithm behavior, and the default values used in our experiments. We also note any deviations from the default values used to obtain the results reported in the main text. In App. H, we analyze the computational complexity of our method; we also describe practical implementation choices used to make the method efficient in practice.
@@ Appendix A Ablation Studies
Table 3: GeoBPE (1%) is trained with 1% of PDBs available during pretraining; GeoBPE (task-specific) directly uses PDBs in each downstream task. All use default value parameters in App. I.
@@
GeoBPE (1%) 59.98 90.17 95.00 95.89 87.73 66.28 88.87 94.95 95.95 71.75 84.56 56.37 72.87 63.83 77.55 80.12
GeoBPE 59.19 91.31 94.94 95.94 87.73 66.21 88.65 95.01 95.90 71.96 84.84 56.44 72.98 64.78 77.06 79.93
GeoBPE (task-specific) 60.16 89.93 95.05 95.92 87.73 66.28 88.82 94.98 95.90 71.85 85.92 56.33 72.72 64.78 77.04 80.23
@@
GeoBPE (1%) 40.42 47.55 34.74 32.21 56.78 55.32 44.50 21.65 50.25 84.87 52.26
GeoBPE 40.89 47.17 37.28 35.61 56.65 53.98 45.26 23.60 47.28 85.75 52.21
GeoBPE (task-specific) 39.39 44.00 37.94 38.36 56.22 54.22 45.02 24.22 46.58 84.57 51.79
GeoBPE is task-agnostic, and using task-specific data does not increase performance. For each task $i$, let $\mathsf{T}^{\text{task}}_{i}$ be a tokenizer fitted using only $\mathcal{D}^{\text{train}}_{i}$ (with its own vocabulary $\mathcal{V}_{i}$ but the same feature extractor $F_{\Theta}$), and define $r_{\mathcal{V}_{i}}(\mathbf{x})=\Psi_{\mathcal{V}_{i}}(F_{\Theta}(\mathbf{x}))$. We follow the same downstream transfer evaluation. We find an interesting result in Table 3, where directly training on the task-specific dataset does not meaningful change downstream prediction results. A closer look reveals the underlying reason is because the individual tokens do not differ significantly; motifs added to $\mathcal{V}$, in order, are similar across both GeoBPE and GeoBPE (task-specific). We can interpret this both positively and negatively. GeoBPE is insensitive to task-specific data and learns the “language” of protein folds consistently. This may be desirable for reusability of a tokenizer, as one does not need to retrain it for different data distributions, as all protein folds obey the same universal principles (Petsko & Ringe, 2004). At the same time, this upper bound tests whether the tokenizer can tailor its vocabulary to individual datasets for potentially higher scores, indicating GeoBPE by itself may lack the parameter capacity to overfit to individual tasks.
@@
GeoBPE reaches comparable downstream transfer performance with 1% of data. In Table 3, we see GeoBPE fitted on just $1\%$ of the pretraining data is enough to transfer, on average, the same amount of performance downstream as GeoBPE trained on the full dataset. There are no meaningful differences between GeoBPE and GeoBPE (1%), with GeoBPE doing 1.7% better on physicochemical property prediction and GeoBPE (1%) doing better 0.2% better on functionals ite prediction. These findings can be interpreted both positively and negatively for GeoBPE induced features: (1) they are extremely data-efficient, learning useful signals to transfer downstream with as few as 300 PDB structures; (2) they underfit the data, with no noticeable improvements for more data. Taken together, these findings imply GeoBPE is a lightweight add-on on top of any pretrained features $\Theta$, but feeding more data to GeoBPE yields diminishing returns quickly.
@@
GeoBPE is multi-resolution, revealing finer details as more tokens are introduced. In Fig. 4, we run a coarse-grained version of GeoBPE (small initial $|\mathcal{V}|$) to observe an interesting feature of GeoBPE’s design. As newly introduced tokens re-quantize the occurrences from the original data (span gathering step in Alg. 9, tokenization can adaptively increase the resolution if the new prototypes better capture the modes of variability for those occurrences than their previous quantization. We expose this via hyperparameters bins & num_p (see App. I), which tradeoff the super-resolution effect against coarse-graining effect at different token sizes, offering fine-grained control.

### v2

@@ 3.3 Principles of Protein Structure Tokenization
Principle 3: Downstream transfer via codebook/vocabulary. Let $\mathcal{V}$ be the vocabulary of $\mathsf{T}$ and let $N(\mathbf{x})$ be the residue count. Let $\Theta$ parameterize a pretrained feature extractor. The codebook/vocabulary $\mathcal{V}$ induces per-residue features $r_{\mathcal{V}}(\mathbf{x})=\Psi_{\mathcal{V}}(F_{\Theta}(\mathbf{x}))\in(\mathbb{R}^{d})^{N(\mathbf{x})}$. An ideal tokenizer of protein structures should go beyond pure compression; it should learn useful signals related to function. We loosely define the ability of a PST to transfer useful signals by test performance on a battery of downstream tasks when parameterizing samples $\mathbf{x}$ by the vocabulary $\mathcal{V}$ together with a feature extractor $\Theta$. We benchmark downstream transfer of GeoBPE against others in Table 1 (GeoBPE-Transfer).
@@ 4 Experiments
  • Token Efficiency: (Q5) How frequent and balanced is vocabulary utilization? (Q6) Does small-scale language modeling generate better structures with GeoBPE or VQ-VAE tokens?
@@
Downstream Transfer. For VQ-VAEs, $\Theta$ and $\mathcal{V}$ are jointly learned, so we set $r^{\text{VQ-VAE}}_{\mathcal{V}}(\textbf{x}):=r^{\text{VQ-VAE}}(\textbf{x})\leftarrow\text{Enc}(\textbf{x})$. For GeoBPE-Transfer, we use $\Theta\leftarrow\text{ESM3}$ to demonstrate how $\mathcal{V}^{\text{{GeoBPE}}}$ can transfer useful signals from $\mathcal{F}_{\Theta}(\textbf{x})$ to $r_{\mathcal{V}}(\mathcal{F}_{\Theta}(\textbf{x}))$.
Performance Metrics. Compression measures Bits-Per-Residue (BPR), as defined in Sec. 3.3. Distortion ($\Delta$) use standard RMSD and LDDT. Token Efficiency uses Codebook Utility Rate (UR), Perplexity (details in App. D) and Small Structure Language Model Evaluation (SSLM-Eval) (details in App. E). SSLM-Eval compares tokenizers (GeoBPE vs VQ-VAEs) via integration with a small $\sim$7.3M Transformer architecture after respectively tokenizing the pretraining data splits (Algo. 2). Under the same data, model, training and compute resources, the respective models generate new sentences, detokenizes them into structures, and we compare relative generation metrics (Algo. 3, 4, 5). Downstream Transfer covers 12 tasks (24 test splits) using AUROC (%) for functional site prediction, Spearman’s $\rho$ (%) for flexibility prediction, and Macro F1 (%) for fold classification. Expert Agreement measures Domain & Segment Recall/Precision/F1/IOU (details in App. G).
Computational Complexity / Implementation Details. We analyze the theoretical complexity of GeoBPE in App. J and justify the steps we took towards efficient implementation and use.
@@ 5 Results
Figure 4: Plots of $(\operatorname{BPR}(\mathsf{T};\mathcal{D}),\ \Delta(\mathsf{T};\mathcal{D}_{\mathrm{test}}))$ across tokenizers for $\Delta\in\{\text{RMSD},\text{LDDT}\}$. We vary $|\mathcal{V}|\in\{128,256,512,1024\}$ for VQ-VAE/AminoASeed and $|\mathcal{V}|\in\{600,2500,6000,21000\}$ for GeoBPE to sample multiple points; we observe GeoBPE sweeps a smooth tradeoff curve. Hyperparameters in App. K.
@@
Token Efficiency. We report UR & Perplexity averaged over held-out test sets to gauge codebook/vocabulary usage on unseen data, the setting where the tokenizer is deployed. In Table 10, we see all methods except VQ-VAE and ESM3 achieve an average UR of $40\%$; all except VQ-VAE achieve $0.2$ average Perplexity. An ideal tokenizer avoids codebook collapse, but exactly uniform token usage may not be desirable. We introduce SSLM-Eval to stress test whether codebook efficiency actually leads to generative efficiency. SSLM-Eval is a holistic way to compare tokenizers using both encoder token efficiency and decoder’s generative efficiency. In Table 11, we find GeoBPE-Transfer is capable of generating $99\%$ unique and designable backbones, achieving up to $49\%$ higher scTM and maintaining higher diversity than both VQ-VAE methods using the same data splits. We visualize some realistic, novel backbones GeoBPE-Transfer generated in App. E.4. Interestingly, the “less-efficient” VQ-VAE generated $58\%$ more diverse backbones, demonstrating uniform token usage can be counterproductive to language modeling.
Downstream Task Transfer. In Table 1, we see GeoBPE-induced features rank first, on average, across both function and structure property prediction tasks. The relative performance gaps $15.44\%$ and $43.28\%$ quantify the add-on benefits of GeoBPE-induced features. GeoBPE-induced features reverse the trend that discrete PSTs produce less informative representations for downstream tasks (due to quantization-related issues (Yuan et al., 2025)), highlighting that hierarchical structure from discrete vocabularies raises the ceiling on downstream transfer.
Further Ablations. We include a comprehensive series of ablation studies in App. A demonstrating GeoBPE’s data-efficiency, GeoBPE-Transfer’s task-agnosticism, GeoBPE tokenizer’s scalability, GeoBPE tokens’ adaptive resolution over iterations, and performance vs runtime tradeoffs in components (1), (3) & (4). Key findings include: (i) GeoBPE shows better OOD generalization when fitted on 1% training data; (ii) GeoBPE-Transfer predictions are no worse when GeoBPE was fitted with (a) 1% of the pretraining PDBs, (b) the downstream task-specific PDBs; (iii) GeoBPE performance gains diminish beyond $M_{\max}=5000$ randomly sampled motif occurrences used to extract prototypes, taming a complexity term that depends on $M_{\max}$.
Table 1: Downstream transfer performance benchmark. We underline and bold the best continuous and discrete PSTs, respectively;   indicates the best method across both. The relative performance v.s. ESM3 for GeoBPE-Transfer is included. Omitted rows in Table 6; GeoBPE hyperparameters are in App. K.
@@
ProteinMPNN MIF FoldSeek ProTokens ESM3 VanillaVQ AminoAseed GeoBPE-Transfer (v.s. ESM3)
@@
BindInt Fold 51.83 50.38 53.18 44.66 44.30 47.25 47.11 59.19 (+33.61%)
SupFam 94.00 94.56 46.26 86.05 90.77 86.71 90.53 91.31 (+0.59%)
BindBio Fold 78.42 85.79 32.37 58.47 62.84 62.02 65.73 94.94 (+51.08%)
SupFam 81.00 87.27 52.44 60.47 65.22 62.92 68.30 95.94 (+47.10%)
BindShake Org 75.52 79.90 53.43 59.10 66.10 67.04 69.61 87.73 (+32.72%)
CatInt Fold 61.05 59.62 53.43 58.16 61.09 58.89 62.19 66.21 (+8.38%)
SupFam 93.40 96.49 51.41 83.85 89.82 85.00 91.91 88.65 (-1.30%)
CatBio Fold 82.49 85.85 56.33 67.68 65.33 67.58 65.95 95.01 (+45.43%)
SupFam 93.19 96.97 53.78 64.05 74.65 70.92 87.59 95.90 +28.47%
Con Fold 57.18 58.43 49.20 57.20 55.22 56.98 57.23 71.96 (+30.32%)
SupFam 84.68 92.66 51.31 70.64 80.53 74.60 86.60 84.84 (+5.35%)
@@ 6 Discussion
Table 2: We annotate 100 PDBs from each dataset and report % of 1,000 random equal-length segmentations that GeoBPE matches or outscores. Omitted columns are in Table 7. Secondary structure analysis in Table 15.
@@
We ran CATH Functional Families (FunFams) (Das et al., 2015b) to obtain domain boundaries and compared them against GeoBPE-derived motifs. Because sequence conservation is linked to structural preservation, we expect overlap between predicted motifs and functional domains. In Table 2, GeoBPE achieves 99.97% domain recall with mean F1 = 0.996 and IOU = 0.992, showing near-perfect agreement across 10 datasets. The agreement is not only geometric but also functional: GeoBPE tokens frequently coincide with boundaries of ligand-binding grooves, transmembrane cavities, and scaffolding helices, capturing motifs that underlie molecular recognition and catalysis. This suggests GeoBPE does more than segment folds consistently: it surfaces interpretable structural primitives that map onto biochemical roles, offering a functional vocabulary absent in prior PSTs. Details are in App. G
Case Study: Human Expert Analysis of Interpretability. We conducted three expert evaluations of GeoBPE-derived hierarchies (App. H). Across proteins, the discovered motifs align with functionally meaningful substructures, including regions mediating ligand binding, molecular recognition, and structural gating. In the SLC25A20 transporter (Fig. 9), GeoBPE isolates a transmembrane binding cavity formed by helices and polar residues. In the 14-3-3:Tau complex (Fig. 10), it identifies a canonical phospho-binding groove stabilized by charged side chains. Recurrent local motifs (aromatic cages, polar bridges, helix-helix clamps) are combined into higher-order scaffolds that mirror established biochemical organization. These hierarchies capture geometric regularities and also modular design principles conserved across folds and families. Even in compact domains, such as nucleotide-recognition modules, GeoBPE motifs reveal the coupling between geometric curvature and chemical specificity, meaning that GeoBPE surfaces reusable motifs that are both interpretable and evolutionarily grounded.
Limitations. GeoBPE currently does not incorporate sequence or side chains, but can via direct extensions, e.g. taking the Cartesian product of the current vocabulary with amino acid types and augmenting the backbone formulation with type-dependent $\chi$-angle spans. The present integration with small-scale Transformers is set up to compare tokenizers’ compability with language modeling on structure tokens; only relative backbone design metrics are relevant. Generative performance depends on model capacity and data scale, which are orthogonal to the tokenizer. In the separate SSLM scalability study, we see steep gains in generative performance when both LM parameter count and pretraining data increase ten-fold, conforming with scaling law expectations. The improved numbers are preliminary evidence of GeoBPE’s promise as a tokenizer for lage-scale PLMs, but are not competitive with state-of-the-art backbone design models.
@@ 7 Conclusion
Reproducibility Statement
Code for GeoBPE and steps to reproduce all experiments in the paper are available at https://github.com/shiningsunnyday/PT-BPE/. We include detailed descriptions for understanding our method in the main text, with mathematical descriptions and pseudocodes in the Appendix. In App. K, we list the key hyperparameters, their effects on algorithm behavior, the default values used in our experiments, and any deviations from the default values used to obtain the results reported in the main text. In App. J, we analyze the computational complexity of our method and describe practical implementation choices used to make the method efficient in practice.
Acknowledgments
M.S. and M.Z. gratefully acknowledge partial support by NSF CAREER Award 2339524, ARPA-H Biomedical Data Fabric (BDF) Toolbox Program, Amazon Faculty Research, Google Research Scholar Program, AstraZeneca Research, GlaxoSmithKline Award, Roche Alliance with Distinguished Scientists (ROADS) Program, Sanofi iDEA-iTECH Award, Boehringer Ingelheim Award, Merck Award, Optum AI Research Collaboration Award, Pfizer Research, Gates Foundation (INV-079038), Chan Zuckerberg Initiative, Collaborative Center for XDP at Massachusetts General Hospital, John and Virginia Kaneb Fellowship at Harvard Medical School, Biswas Computational Biology Initiative in partnership with the Milken Institute, Harvard Medical School Dean’s Innovation Fund for the Use of Artificial Intelligence, and the Kempner Institute for the Study of Natural and Artificial Intelligence at Harvard University. Any findings, conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the funders.
@@ Appendix A Ablation Studies
GeoBPE is task-agnostic, and using task-specific data does not increase downstream perfor- mance of GeoBPE-Transfer. For each task $i$, let $\mathsf{T}^{\text{task}}_{i}$ be a tokenizer fitted using only $\mathcal{D}^{\text{train}}_{i}$ (with its own vocabulary $\mathcal{V}_{i}$ but the same feature extractor $F_{\Theta}$), and define $r_{\mathcal{V}_{i}}(\mathbf{x})=\Psi_{\mathcal{V}_{i}}(F_{\Theta}(\mathbf{x}))$. We follow the same downstream transfer evaluation. We find an interesting result in Table 3, where directly training on the task-specific dataset does not meaningful change downstream prediction results. A closer look reveals the underlying reason is because the individual tokens do not differ significantly; motifs added to $\mathcal{V}$, in order, are similar across both GeoBPE-Transfer and GeoBPE-Transfer (task-specific). We can interpret this both positively and negatively. GeoBPE is insensitive to task-specific data and learns the “language” of protein folds consistently. This may be desirable for reusability of a tokenizer, as one does not need to retrain it for different data distributions, as all protein folds obey the same universal principles (Petsko and Ringe, 2004). At the same time, this upper bound tests whether the tokenizer can tailor its vocabulary to individual datasets for potentially higher scores, indicating GeoBPE by itself may lack the parameter capacity to overfit to individual tasks.
Table 3: GeoBPE-Transfer (1%) runs Algo. 1 with 1% of the pretrain training set, then uses the output vocabulary to induce features; GeoBPE-Transfer (task-specific) does not use pretraining data; instead it runs Algo. 1 with downstream data to learn a vocabulary. All use default value parameters in App. K.
@@
GeoBPE-Transfer (1%) 59.98 90.17 95.00 95.89 87.73 66.28 88.87 94.95 95.95 71.75 84.56 56.37 72.87 63.83 77.55 80.12
GeoBPE-Transfer 59.19 91.31 94.94 95.94 87.73 66.21 88.65 95.01 95.90 71.96 84.84 56.44 72.98 64.78 77.06 80.20
GeoBPE-Transfer (task-specific) 60.16 89.93 95.05 95.92 87.73 66.28 88.82 94.98 95.90 71.85 85.92 56.33 72.72 64.78 77.04 80.23
@@
GeoBPE-Transfer (1%) 40.42 47.55 34.74 32.21 56.78 55.32 44.50 21.65 50.25 84.87 52.26
GeoBPE-Transfer 40.89 47.17 37.28 35.61 56.65 53.98 45.26 23.60 47.28 85.75 52.21
GeoBPE-Transfer (task-specific) 39.39 44.00 37.94 38.36 56.22 54.22 45.02 24.22 46.58 84.57 51.79
@@
GeoBPE-Transfer maintains comparable downstream transfer performance even when GeoBPE was fitted on 1% of pretraining data. In Table 3, we see GeoBPE fitted on just $1\%$ of the pretraining data is enough to transfer, on average, the same amount of performance downstream as GeoBPEfitted on the full dataset. There are no meaningful differences between GeoBPE-Transfer and GeoBPE-Transfer (1%), with GeoBPE-Transfer doing 1.7% better on physicochemical property prediction and GeoBPE-Transfer (1%) doing better 0.2% better on functionals ite prediction. These findings can be interpreted both positively and negatively for GeoBPElearned vocabularies: (1) they are extremely informative, learning useful signals to transfer downstream with as few as 300 PDB structures; (2) they underfit the data, with no noticeable improvements from using more data to learn the tokenizer. Taken together, these findings imply GeoBPE is a lightweight add-on on top of any pretrained features $\Theta$, but feeding more data to GeoBPE yields diminishing downstream returns quickly.
Table 4: We use GeoBPE-Transfer from Table 1 (reported as 100%), then vary only the percent of downst- ream task training data available, fixing the same valid and test sets.
Task Split GeoBPE-Transfer
20% 40% 60% 80% 100%
Functional Site Prediction (AUROC%)
BindInt Fold 58.78 61.18 59.79 59.45 59.19
SupFam 89.55 89.88 90.42 90.71 91.31
BindBio Fold 94.86 94.82 94.96 94.97 94.94
SupFam 95.69 95.79 95.88 95.97 95.94
BindShake Org 87.40 87.59 87.65 87.64 87.73
CatInt Fold 64.66 65.55 66.94 66.29 66.21
SupFam 87.96 88.27 88.78 88.73 88.65
CatBio Fold 94.95 94.92 94.94 94.96 95.01
SupFam 95.67 95.81 95.96 95.97 95.90
Con Fold 71.42 71.37 71.72 71.74 71.96
SupFam 83.77 84.02 84.76 84.76 84.84
Rep Fold 55.04 54.25 56.02 56.41 56.44
SupFam 73.24 75.18 75.89 71.58 72.98
Ept Fold 63.63 53.28 61.72 61.66 64.78
SupFam 71.21 49.39 73.59 76.01 77.06
Average AUROC% 79.19 77.42 79.93 79.79 80.20
Physicochemical Property Prediction (Spearman’s $\rho$%)
FlexRMSF Fold 41.49 43.33 38.48 39.02 40.89
SupFam 34.74 45.41 44.61 47.06 47.17
FlexBFactor Fold 23.90 27.86 33.83 34.82 37.28
SupFam 23.80 25.46 37.70 36.49 35.61
FlexNEQ Fold 54.40 56.07 55.98 57.53 56.65
SupFam 51.12 53.77 52.52 54.96 53.98
Average $\rho$% 38.24 41.98 43.85 44.98 45.26
Structure Property Prediction (Macro F1%)
Homo Fold 14.35 20.55 23.74 24.25 23.60
SupFam 27.63 35.11 43.09 43.98 47.28
Fam 65.79 73.66 80.77 82.04 85.75
Average Macro F1% 35.92 43.11 49.20 50.09 52.21
GeoBPE-Transfer does not underfit the data for structure-related downstream tasks. In Table 4, we see for physicochemical (residue-level regression) and fold-level tasks, the model indicates a clear propensity for more training data, with step-wise gains for every 20% of training data. $20\%\rightarrow 100\%$ training data sees performance lift significantly ($18.89\%$ average $\rho\%$ and $45.35\%$, respectively). For residue-level classification tasks, the lift is only marginal ($1.28\%$). We hypothesize the cause is not limited capacity, but rather that the tasks are localized label predictions; a residue-level receptive field is sufficient when features are informative. This shows GeoBPE’s data-efficiency at learning a vocabulary does not limit its capacity on downstream tasks that require multi-scale resolution (e.g. structural flexibility or fold-level classification), which existing fixed-size tokenizers cannot at both residue and structure-level. Thus, the data-efficiency strengths of GeoBPE training is orthogonal to downstream modeling. The structure-related tasks in Table 4 see large gains in performance with more training data, implying GeoBPE-Transfer scales to complex, hierarchical structural patterns that can only be learned from more data.
Table 5: We rerun the GeoBPE experiments used to trace out the Pareto Front in Fig. 4 by using $1\%$ of pretraining data. We include raw numbers of Fig. 4 (bottom rows) for comparison. All other hyperparameter settings are kept the same (App. K).
Train Valid CAMEO CASP14
RMSD LDDT RMSD LDDT RMSD LDDT RMSD LDDT
GeoBPE (1%) ($|V|=600$) 1.72 0.74 1.63 0.73 1.66 0.73 1.53 0.72
GeoBPE          ($|V|=600$) 1.66 0.73 1.71 0.72 1.77 0.72 1.53 0.72
GeoBPE (1%) ($|V|=2278$) 1.57 0.75 1.51 0.71 1.51 0.74 1.43 0.73
GeoBPE          ($|V|=2500$) 1.41 0.75 1.50 0.74 1.57 0.74 1.51 0.73
GeoBPE (1%) ($|V|=5278$) 1.36 0.77 1.34 0.76 1.35 0.75 1.30 0.74
GeoBPE          ($|V|=6000$) 1.37 0.76 1.46 0.75 1.52 0.74 1.54 0.72
GeoBPE (1%) ($|V|=20278$) 1.29 0.77 1.28 0.76 1.28 0.76 1.37 0.73
GeoBPE          ($|V|=21000$) 1.21 0.77 1.28 0.76 1.40 0.75 1.55 0.72
GeoBPE is data-efficient OOD, but more training data can lower training distortion. In Table 5, we see GeoBPE (1%) consistently achieves lower distortion ($\downarrow 7.7\%$ RMSD averaged, $\uparrow 0.06$ LDDT summed across all four runs and both test splits) than GeoBPE. This suggests a small, well-chosen set of structures is enough for GeoBPE to achieve superior reconstruction on OOD structures, and more training data can introduce noise and hinder generalization. However, GeoBPE (1%) obtains $\uparrow 5.1\%$ averaged, $\downarrow 0.02$ LDDT summed across all four runs on the respective training splits. A lower RMSD suggests GeoBPE better captures global structural fidelity; by constructing the vocabulary from the full pretraining dataset, it can choose more representative prototypes; hence, its vocabulary better preserves global fold. Meanwhile, a slightly lower LDDT indicates GeoBPE (1%) can capture a few local details in the 1% subset of structures better than GeoBPE. This suggests GeoBPE (1%) is more sensitive to the individual local interactions of the small set of structures it fitted with; GeoBPE considers vastly more structures. In summary, GeoBPE is preferred for fold-preserving compression of whole datasets, but GeoBPE (1%) can be feasible if not superior when GeoBPE is primarily used to tokenize unseen data.
@@
GeoBPE is multi-resolution, revealing finer details as more tokens are introduced. In Fig. 5, we run a coarse-grained version of GeoBPE (small initial $|\mathcal{V}|$) to observe an interesting feature of GeoBPE’s design. As newly introduced tokens re-quantize the occurrences from the original data (span gathering step in Alg. 9, tokenization can adaptively increase the resolution if the new prototypes better capture the modes of variability for those occurrences than their previous quantization. We expose this via hyperparameters bins & num_p (see App. K), which tradeoff the super-resolution effect against coarse-graining effect at different token sizes, offering fine-grained control.